Discuss the maxima and minima of f(x,y)=x3y2(24 –x - y).


1
Expert's answer
2021-02-25T16:23:32-0500

"f(x,y) = x^3y^2(24-x-y) = 24x^3y^2-x^4y^2-x^3y^3"

necessary extremum condition:

Let "M_0(x_0, y_0)" - extremum, then:

"\begin{cases}
\cfrac{\partial f}{\partial x}(x_0,y_0) = 0\\\cfrac{\partial f}{\partial y}(x_0,y_0) = 0
\end{cases}"


Find first partial derivatives:

"\cfrac{\partial f}{\partial x} = 72x^2y^2 - 4x^3y^2-3x^2y^3 \\
\cfrac{\partial f}{\partial y} = \\
 \begin{cases} 
 \cfrac{\partial f}{\partial x} = 72x^2y^2 - 4x^3y^2-3x^2y^3 = 0 \\
 \cfrac{\partial f}{\partial y} = 48x^3y-2x^4y-3x^3y^2=0 \\
 \end{cases} \\"

Solve this system for "x,y":

"\begin{cases}
x^2y^2(72-4x-3y) = 0\\
x^2y(48x-2x^2-3y) = 0
\end{cases}\\
\begin{cases}
x^2y^2 = 0 \,or\,-4x-3y+72 = 0\\
x^2y =0\, or\,-2x^2+48x-3y=0
\end{cases}"

Find the roots of first line:

"x^2y^2 = 0 \,or\,-4x-3y+72 = 0"

From second equation:

"y = 24 - \cfrac{4}{3}x"

And put it in first equation:

"24x-\cfrac{4}{3}x^2 = 0\\
-4x^2+72x=0\\
x(-4x+72) =0\\
x=0\;or \; x =18\\
y=24\; or\;y=0"

"M_1(0;24),M_2(18;0)"

Find the roots of second line:

"x^2y=0\, or\,(48x-2x^2-3y) = 0"

From second equation:

"y = -\cfrac{2}{3}x^2+16x"

And put it in first equation:

"x^2(-\cfrac{2}{3}x^2+16x)=0\\
x^3(-\cfrac{2}{3}x+16)= 0\\
x = 0\;or\; x =24\\
y=0\; or \;y=0\\
M_3(0;0), M_4(24;0)"


"\begin{cases}
M_1(0;24),M_2(18;0)\\
M_3(0;0), M_4(24;0)
\end{cases}"

The obtained points are the solution of two  aggregates of equations, but the necessary extremum condition says that the points must be a solution to the system. Therefore, the system has one solution - "M_3(0;0)" . So we can check sufficient condition for extremum:

If "\cfrac{\partial^2 f}{\partial x^2}(x_0;y_0) = A, \cfrac{\partial^2 f}{\partial y^2}(x_0;y_0) = B, \cfrac{\partial^2 f}{\partial x \partial y}(x_0;y_0) = C" ,

then :

if "AC - B^2 > 0\, and\, A>0 \,(A<0), then\, M(x_0;y_0) - minimum (maximum)"

If "AC-B^2 < 0, then\, M(x_0;y_0) -isn't\; extremum"

if "AC-B^2 = 0, \; then \, we \,can't\, say \,anything."


"\cfrac{\partial^2 f}{\partial x^2} = 144xy^2 -12x^2y^2-6xy^3\\
\cfrac{\partial^2 f}{\partial y^2} = 48x^3-2x^4-6x^3y\\
\cfrac{\partial^2 f}{\partial x \partial y} = 144x^2y - 8x^3y-49x^2y^2"

"A(0;0) = 0\;B(0;0) = 0, \;C(0;0) = 0"

"AC-B^2 = 0", so the point "M_3(0;0)" is not a minimum or maximum point.